# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

### v3

@@ 8 Discussion
In conclusion, we introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in multi-label classification. Our experiments, particularly with the FairPO-CPO variant, highlight the value of nuanced preference signals for navigating complex fairness-performance trade-offs in challenging discrimination tasks. While promising, FairPO has technical limitations: its dynamic confusing set can lead to instability or sparse signals for rare labels; the DPO-based variant and non-privileged constraint rely on a well-calibrated reference model; and GRPO’s balancing of heterogeneous losses requires careful tuning and lacks theoretical convergence guarantees. These challenges also define our future work, which includes extending FairPO’s principles to multi-label attribute generation (Appendix A), conducting comprehensive empirical validation on more datasets and modalities, analyzing alternative label partitioning strategies, and pursuing theoretical insights into the framework’s convergence properties.
Appendix A Adapting FairPO for Multi-Attribute Generation
@@
Planned Experiments: The planned task is generating text from a prompt, requiring preference datasets where the driving attribute $j_{i}$ is known. This could involve partitioning attributes like ‘Helpfulness’ or ‘Harmlessness’ from existing RLHF datasets (e.g., Anthropic HH-RLHF [Bai and Others, 2022], Summarization preferences [Stiennon et al., 2022]) or creating new, explicitly annotated data. We plan to fine-tune pretrained language models (e.g., Gemma [Team, 2024], Llama [Touvron and Others, 2023], Mistral [Jiang and Others, 2023]) with this objective. Evaluation would rely on preference-based metrics, such as win rates for privileged and non-privileged attribute groups, compared against baselines like standard DPO and SFT. Future experiments would also explore hyperparameter effects and adapting the SimPO/CPO loss formulations for generation.
Appendix B FairPO Algorithm
@@
This entire process repeats for a predefined number of iterations or until convergence, allowing FairPO to dynamically balance its objectives to achieve robust fairness. For variants like FairPO-SimPO or FairPO-CPO, the core logic remains identical; only the DPO-inspired preference loss component is replaced with their respective preference formulations (e.g., Eq. 12 or 13). The overall GRPO structure and non-privileged handling are consistent across all variants.
@@
Appendix C Dataset and Preprocessing Details
MS-COCO 2014 [Lin et al., 2015]: We used the official 2014 train/val splits. The training set contains 82,783 images and the validation set (used as our test set) contains 40,504 images. There are 80 object categories. The privileged set $\mathcal{P}$ consisted of the 16 labels (20% of 80) with the lowest frequency in the training set. The remaining 64 labels formed $\bar{\mathcal{P}}$.
NUS-WIDE [Chua et al., 2009]: This dataset contains 269,648 images with 81 concept labels. We used the common split of 161,789 images for training and 107,859 for testing. The privileged set $\mathcal{P}$ consisted of the 16 labels (approx. 20% of 81) with the lowest frequency in the training set. The remaining 65 labels formed $\bar{\mathcal{P}}$.
@@ Appendix C Dataset and Preprocessing Details
Appendix D Baseline Experimental Details
This section provides further details on the implementation and hyperparameter tuning for the baseline methods used in our experiments. Unless otherwise specified, all baselines were trained using the same base Vision Transformer (ViT) architecture, optimizer (AdamW), number of epochs, and batch size as the main FairPO experiments for fair comparison. Hyperparameters specific to each baseline were tuned on the validation set of MS-COCO and NUS-WIDE. The privileged group $\mathcal{P}$ comprised the 20% least frequent labels, and the non-privileged group $\bar{\mathcal{P}}$ comprised the remaining 80%.
D.1 BCE-SFT (Reference Model)
The BCE-SFT baseline represents standard supervised fine-tuning. Its objective is to minimize the sum of independent Binary Cross-Entropy (BCE) losses for each of the $T$ labels, where the loss for an instance $(x_{i},y_{i})$ is $\mathcal{L}_{\text{BCE-SFT}}=\sum_{t=1}^{T}\text{BCE}(m(x_{i};\mathbf{w}_{t}),y_{it})$. For implementation, a separate non-linear MLP classifier head is trained for each label $t$ on top of frozen ViT features, consistent with the main FairPO architecture. The primary hyperparameter tuned was the learning rate for these classifier heads, selected from $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$ to achieve the best overall mean Average Precision (mAP) on the validation set. Weight decay was set to 0.01, matching the FairPO experiments.
This model serves two critical roles in our study. First, it acts as a direct performance baseline against which we measure fairness improvements. Second, the final trained parameters $\{\mathbf{w}_{t}\}$ from this SFT process become the reference parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}=\{\hat{\mathbf{w}}_{t}\}$ used throughout the FairPO framework, both for the DPO-based losses and for the non-privileged loss constraint. The model was trained for the same number of epochs as FairPO, with early stopping based on overall validation mAP.
D.2 BCE-SFT + Privileged Re-Weighting (RW)
This baseline modifies the standard BCE-SFT by applying static loss re-weighting to improve performance on the privileged labels $\mathcal{P}$. The objective is to assign a higher weight to their contribution in the total loss function:
$\mathcal{L}_{\text{RW}}=\lambda_{\mathcal{P}}\sum_{l\in\mathcal{P}}\text{BCE}(m(x_{i};\mathbf{w}_{l}),y_{il})+\lambda_{\bar{\mathcal{P}}}\sum_{j\in\bar{\mathcal{P}}}\text{BCE}(m(x_{i};\mathbf{w}_{j}),y_{ij})$
For this setup, we use the same per-label classifier architecture as BCE-SFT, but the loss computation is modified to incorporate the weights. The key hyperparameter is the privileged group weight $\lambda_{\mathcal{P}}$, which was tuned via grid search from values $\{2,3,5,8,10\}$. The value was chosen to maximize mAP on the privileged group without an excessive drop in performance on the non-privileged group. For our reported results, a weight of $\lambda_{\mathcal{P}}=5$ was found to be effective, while $\lambda_{\bar{\mathcal{P}}}$ was fixed at $1$. Training was conducted similarly to BCE-SFT.
D.3 Group DRO + BCE
This baseline applies the Group Distributionally Robust Optimization technique to the standard BCE loss. The objective is to minimize the worst-case BCE loss across the predefined privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) groups:
$\min_{\{\mathbf{w}_{t}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\geq 0,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}=1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\text{BCE}}(\mathcal{P})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\text{BCE}}(\bar{\mathcal{P}})\right]$
Here, $\mathcal{L}_{\text{BCE}}(\mathcal{G})$ is the average BCE loss for all labels within a group $\mathcal{G}$. The implementation uses the same BCE-SFT architecture. Per-label BCE losses are computed and then aggregated separately for the $\mathcal{P}$ and $\bar{\mathcal{P}}$ groups. Model parameters $\{\mathbf{w}_{t}\}$ are updated with AdamW, while the group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are updated iteratively using a mirror ascent step: $\alpha_{\mathcal{G}}^{(s+1)}\propto\alpha_{\mathcal{G}}^{(s)}\exp(\eta_{\alpha}\cdot\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{G}))$. The final loss for backpropagation is the adaptively weighted sum of the group losses.
For hyperparameters, the model learning rate was tuned as for BCE-SFT. The group weight learning rate $\eta_{\alpha}$ was tuned from $\{0.01,0.05,0.1,0.2\}$, with $\eta_{\alpha}=0.05$ selected for its stable convergence and good worst-group performance. For the reported results, raw group losses were used in the exponential update, though we note that a loss scaling approach similar to that in FairPO could also be employed for enhanced stability. The training protocol regarding epochs and batching was identical to FairPO.
Appendix E FairPO Experimental Details
E.1 Common Setup for All FairPO Variants

### v4

@@ 7 Discussion
In conclusion, we introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in MLC, with our CPO variant proving particularly effective. While promising, FairPO has limitations, including the potential instability of its dynamic ‘confusing set‘, the DPO variant’s reliance on a reference model, and the need for careful tuning of GRPO’s balancing act. These challenges directly motivate our future work, which will focus on extending FairPO to attribute generation (Appendix D), performing broader empirical validation, exploring alternative label partitioning strategies, and developing theoretical insights into the framework’s convergence.
Appendix A Preliminaries: Preference Optimization Methods
Direct Preference Optimization (DPO): DPO (rafailov2023direct) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$. Assuming a Bradley-Terry preference model tied to an implicit reward function related to $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, DPO maximizes the likelihood of observed preferences, resulting in the loss:
Group Robust Preference Optimization (GRPO): GRPO (ramesh2024group) extends preference optimization to handle diverse preference groups $\{D_{g}\}_{g=1}^{K}$. Instead of minimizing the average loss, GRPO minimizes the worst-case loss across groups using a robust objective:
Simple Preference Optimization (SimPO): SimPO (meng2024simposimplepreferenceoptimization) aims to align the implicit reward with generation metrics and eliminates the need for $\pi_{\text{ref}}$. It uses the length-normalized average log-likelihood as the reward: $r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y|x)$. It also introduces a target margin $\gamma>0$ into the preference model. The resulting SimPO loss is:
Contrastive Preference Optimization (CPO): CPO (xu2024contrastivepreferenceoptimizationpushing) also removes the dependency on $\pi_{\text{ref}}$ for efficiency, approximating the DPO objective. It combines a reference-free preference loss with a negative log-likelihood (NLL) regularizer on preferred responses $y_{w}$ to maintain generation quality:
$L_{\text{CPO}}(\pi_{\theta})$ $=L_{\text{prefer}}+L_{\text{NLL}}.$ (8)
Appendix B FairPO: Detailed Methodology
B.1 Objective for the Privileged Group: Preference-Based Discrimination
For the privileged group $\mathcal{P}$, simply adjusting the weight of a standard BCE loss is insufficient. The primary challenge for these labels is often not basic classification, but fine-grained discrimination against closely related, incorrect alternatives. To address this directly, we reformulate the learning objective for this group as a preference learning task as motivated in our introduction.
Defining Confusing Counterparts.
For a given instance $\mathbf{x}_{i}$ and a privileged label $l\in\mathcal{P}$, we dynamically identify a set of confusing counterparts based on the model’s current predictions. We define two such sets:
  • When the true label is positive ($y_{il}=1$), the set of confusing negatives is composed of incorrect labels that the model scores higher than or equal to the true label:
    $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(\mathbf{x}_{i};\mathbf{w}_{k})\geq m(\mathbf{x}_{i};\mathbf{w}_{l})\}.$ (9)
  • When the true label is negative ($y_{il}=0$), the set of confusing positives is composed of correct labels that the model scores lower than or equal to the true negative:
    $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=1\text{ and }m(\mathbf{x}_{i};\mathbf{w}_{k})\leq m(\mathbf{x}_{i};\mathbf{w}_{l})\}.$ (10)
The overall confusing set for $(i,l)$ is denoted by $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
Conditional Objective with a Stablility Fallback.
Our objective for the privileged group is designed to be adaptive. When a confusing set $S_{il}$ is non-empty, we apply a preference loss $\mathcal{L}_{\text{pref}}$ to directly target these hard discriminative cases. However, as the model learns and becomes more accurate, the confusing set for many examples will naturally become sparse or empty. Relying solely on the preference loss in such a scenario would lead to an unstable, sparse gradient signal, effectively stalling the training process. To ensure continuous and stable learning, we incorporate a crucial stability fallback. If $S_{il}=\emptyset$ for a given instance, we revert to a standard Binary Cross-Entropy (BCE) loss:
$\ell_{\text{BCE}}(\mathbf{x}_{i},y_{il};\mathbf{w}_{l})=-[y_{il}\log m(\mathbf{x}_{i};\mathbf{w}_{l})+(1-y_{il})\log(1-m(\mathbf{x}_{i};\mathbf{w}_{l}))].$ (11)
This fallback mechanism is a critical design choice for robustness. Initially, the model makes many mistakes, leading to large confusing sets and frequent application of the preference loss. As training progresses, the preference loss successfully resolves these hard cases, and the BCE fallback takes over to fine-tune the decision boundaries on the now easier examples. Our analysis of the training dynamics confirms that a sufficient preference signal exists throughout training, with the fallback rate gradually increasing as the model improves. The total loss for the privileged group, $\mathcal{L}_{\mathcal{P}}$, is the expectation over this adaptive, conditional choice. In our framework, we explore several modern preference optimization techniques to instantiate $\mathcal{L}_{\text{pref}}$.
FairPO-DPO Variant.
To define our preference losses consistently, let us denote the score of the preferred label in a pair as $m(\mathbf{x}_{i};\mathbf{w}_{p})$ and the score of the dispreferred label as $m(\mathbf{x}_{i};\mathbf{w}_{d})$. The assignment of these roles depends on the ground truth. For example, if $y_{il}=1$ and $k\in S_{il}^{\text{neg}}$, the true label $l$ is preferred ($p=l$) and the confusing negative $k$ is dispreferred ($d=k$). Conversely, if $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$, the true negative $l$ is preferred ($p=l$) and the confusing positive $k$ is dispreferred ($d=k$).
Inspired by DPO (rafailov2023direct), this variant computes a preference loss relative to a reference model $\hat{\mathbf{w}}$. The preference loss is the negative log-likelihood of the preference, modeled by the difference between the log-probability ratios:
$\mathcal{L}_{\text{pref}}^{\text{DPO}}(\mathbf{x}_{i},p,d)=-\log\sigma\left(\beta\left(\log\frac{m(\mathbf{x}_{i};\mathbf{w}_{p})}{m(\mathbf{x}_{i};\hat{\mathbf{w}}_{p})}-\log\frac{m(\mathbf{x}_{i};\mathbf{w}_{d})}{m(\mathbf{x}_{i};\hat{\mathbf{w}}_{d})}\right)\right).$ (12)
The full DPO-based loss for the privileged group, $\mathcal{L}_{\mathcal{P}}^{\text{DPO}}$, combines this preference loss with the BCE fallback:
$\mathcal{L}_{\mathcal{P}}^{\text{DPO}}=\mathbb{E}_{(\mathbf{x}_{i},l)\text{ s.t. }l\in\mathcal{P}}\left[\mathbf{1}_{S_{il}\neq\emptyset}\cdot\mathcal{L}_{\text{pref}}^{\text{DPO}}(\mathbf{x}_{i},p,d)+\mathbf{1}_{S_{il}=\emptyset}\cdot\ell_{\text{BCE}}(\mathbf{x}_{i},y_{il};\mathbf{w}_{l})\right],$ (13)
where for each sampled privileged label $l$ with a non-empty confusing set, a confusing counterpart $k$ is sampled from $S_{il}$, and the pair $(p,d)$ is determined based on $(l,k)$ and the ground truth to compute the loss.
FairPO-SimPO Variant (Reference-Free with Margin).
This variant adapts SimPO (meng2024simposimplepreferenceoptimization) to create a reference-free preference loss that incorporates an explicit target margin $\gamma>0$. We adapt SimPO’s core concept, which was originally designed for sequences, to our per-label score setting. The resulting preference loss is:
$\mathcal{L}_{\text{pref}}^{\text{SimPO}}(\mathbf{x}_{i},p,d)=-\log\sigma\left(\beta\log\frac{m(\mathbf{x}_{i};\mathbf{w}_{p})}{m(\mathbf{x}_{i};\mathbf{w}_{d})}-\gamma\right).$ (14)
The term $\beta\log(\cdot)$ captures the relative preference between the scores, while the margin term $\gamma$ enforces a stronger separation, pushing the model to not just rank the labels correctly but to do so by a significant amount. This preference loss is used in place of the DPO term within the overall conditional objective for the privileged group (analogous to Eq. 13).
FairPO-CPO Variant (Reference-Free with BCE Regularization).
Our final and most effective variant adapts CPO (xu2024contrastivepreferenceoptimizationpushing). This approach is unique in that it integrates a BCE regularizer directly into the preference objective to ensure model stability and a sense of absolute correctness. When a confusing set is found, the preference loss is defined as:
$\mathcal{L}_{\text{pref}}^{\text{CPO}}(\mathbf{x}_{i},p,d,l)=-\log\sigma\left(\beta\log\frac{m(\mathbf{x}_{i};\mathbf{w}_{p})}{m(\mathbf{x}_{i};\mathbf{w}_{d})}\right)+\lambda_{\text{CPO}}\cdot\ell_{\text{BCE}}(\mathbf{x}_{i},y_{il};\mathbf{w}_{l}).$ (15)
For each sampled label $l$, the pair $(p,d)$ is determined from a confusing counterpart $k\in S_{il}$ if the set is non-empty. Here, the first term is a margin-free preference objective, while the second term, weighted by a hyperparameter $\lambda_{\text{CPO}}$, is the standard BCE loss for the privileged label $l$. This BCE component is crucial, acting as a regularizer that grounds the model in learning absolute scores. The full privileged loss objective is then analogous to the DPO formulation (Eq. 13): the comprehensive $\mathcal{L}_{\text{pref}}^{\text{CPO}}$ is used when a confusing set is found, and only the standard BCE term (Eq. 11), is used as a fallback otherwise.
Advantages of the Preference-Based Objective.
Our choice to adapt modern preference optimization techniques is motivated by the fundamental limitations of standard point-wise losses like BCE and Focal Loss for fine-grained discrimination. These losses evaluate each label’s score independently, answering the question, “Is the score for label $l$ correct in isolation?". While effective for general classification, this approach provides an indirect and often insufficient signal for resolving confusion between two closely-scored labels (lin2018focallossdenseobject). For example, a model might correctly learn to assign a high score (e.g., 0.8) to a true positive label and a slightly lower score (e.g., 0.7) to a confusing negative. A point-wise loss would provide only a small error signal for this negative case. The preference-based objective in FairPO is fundamentally different. It operates on pairs of labels and directly answers a more relational question: “Is the score for the correct label $l$ decisively higher than the score for its specific confusing competitor kk?". By optimizing the log-ratio of these scores, our framework generates a strong, targeted gradient to explicitly drive their values apart. This provides a far more direct and effective mechanism for resolving the most challenging discriminative cases within the privileged group, a capability that point-wise losses lack by design.
B.2 Objective for the Non-Privileged Group: Constrained Performance
For the non-privileged group $\bar{\mathcal{P}}$, the primary objective is different. While the model dedicates its capacity to improving the challenging privileged labels, we must ensure that this focus does not come at an unacceptable cost to the performance on the remaining labels. Therefore, our goal for this group is not to aggressively maximize performance, but to preserve it by preventing significant degradation relative to a reliable baseline.
To achieve this, we introduce a constrained objective that leverages a pre-trained reference model with parameters $\hat{\mathbf{w}}$. We use the standard BCE loss (Eq. 11), as our objective for a given label $j\in\bar{\mathcal{P}}$. The model is penalized only if the BCE loss of the current model, $\ell_{\text{BCE}}(\mathbf{x}_{i},y_{ij};\mathbf{w}_{j})$, exceeds the loss of the reference model, $\ell_{\text{BCE}}(\mathbf{x}_{i},y_{ij};\hat{\mathbf{w}}_{j})$, by more than a small, predefined slack margin $\epsilon\geq 0$. This is implemented using a hinge loss mechanism:
$\mathcal{L}_{\bar{\mathcal{P}}}=\mathbb{E}_{(\mathbf{x}_{i},j)\text{ s.t. }j\in\bar{\mathcal{P}}}\left[\max(0,\ell_{\text{BCE}}(\mathbf{x}_{i},y_{ij};\mathbf{w}_{j})-\ell_{\text{BCE}}(\mathbf{x}_{i},y_{ij};\hat{\mathbf{w}}_{j})-\epsilon)\right].$ (16)
The gradient for any non-privileged label is zero as long as its performance is good enough (i.e., close to or better than the reference model). A learning signal is generated only when performance on a label $j$ drops below this safety threshold.
B.3 Group Robust Optimization Formulation
Having defined two distinct objectives, a preference-based loss $\mathcal{L}_{\mathcal{P}}$ for the privileged group and a constrained loss $\mathcal{L}_{\bar{\mathcal{P}}}$ for the non-privileged group, the final challenge is to balance them. As these objectives are often in competition, a simple weighted sum is insufficient. Instead, we formulate the overall training objective as a minimax game, adapting the principles of Group Robust Preference Optimization (GRPO) (ramesh2024group). We treat the privileged and non-privileged label sets as two distinct groups. The goal is to train a model, parameterized by $\mathbf{w}$, that is robust to the worst-case distribution of losses across these groups. This is expressed as the following minimax problem:
$\min_{\{\mathbf{w}_{t}\}}\max_{\alpha\in\Delta}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}\right],$ (17)
where $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$ represent the expected losses for their respective groups, and $\alpha=(\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}})$ is a vector of adaptive weights in the probability simplex $\Delta$. This formulation provides a principled way to manage the performance trade-off between the two groups.
Appendix C Extended Related Works
Fairness in Multi-Label Classification: Ensuring fairness in MLC (zhang2021survey; tantithamthavorn2018impact) is complex due to multi-faceted label relationships. Recent efforts address MLC-specific fairness challenges, such as tackling label imbalance impacting tail labels (wu2023longtailed), learning instance and class-level subjective fairness (gao2023learning), or incorporating fairness in dynamic learning settings like class-incremental MLC (masud2023fairnessaware). FairPO contributes by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets, applying distinct fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and managing them via a robustness framework. This targeted approach to enhancing performance for pre-defined critical labels, while safeguarding others, differentiates our work.
Preference Optimization: Preference optimization, especially for aligning LLMs (ouyang2022training; christiano2017deep), has rapidly advanced. Direct Preference Optimization (DPO) (rafailov2023direct) and its reference-free variants like CPO (xu2024contrastivepreferenceoptimizationpushing) and SimPO (meng2024simposimplepreferenceoptimization) optimize policies directly from preference pairs. The field continues to evolve with methods like Identity Preference Optimization (IPO) for stability (azhar2023ipo), Kahneman-Tversky-based optimization (KTO) (ethayarajh2024kto), simple yet effective Rejection Sampling Fine-Tuning (RFT) (dong2023rejection), and ongoing theoretical analyses of these methods (zhao2024whyisdpo). We use preferences not to rank entire outputs, but to specifically differentiate true label scores from those of their confusing positive/negative counterparts within the privileged label set, thereby sharpening the model’s decision boundaries for critical distinctions.
Group Robust Optimization: Distributionally Robust Optimization (DRO) (ben2009robust), particularly Group DRO (sagawa2019distributionally), aims to improve worst-case performance across predefined data groups, enhancing fairness and robustness. This concept sees continued development, for instance, in improving its practicality (zhang2022fromaverage). Group Robust Preference Optimization (GRPO) (ramesh2024group) extended this to LLM preference learning, balancing performance across preference groups. FairPO directly employs GRPO’s adaptive optimization strategy. However, our groups are defined by the label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$), and GRPO balances their distinct, custom-formulated loss objectives. This provides a principled mechanism for robustly managing the specific fairness-performance trade-offs in our MLC context.
Appendix D Adapting FairPO for Multi-Attribute Generation
@@
Appendix E FairPO Algorithm
@@
This entire process repeats for a predefined number of iterations or until convergence, allowing FairPO to dynamically balance its objectives to achieve robust fairness. For variants like FairPO-SimPO or FairPO-CPO, the core logic remains identical; only the DPO-inspired preference loss component is replaced with their respective preference formulations (e.g., Eq. 14 or 15). The overall GRPO structure and non-privileged handling are consistent across all variants.
@@
Appendix F Dataset and Preprocessing Details
MS-COCO 2014 (lin2015microsoftcococommonobjects): We used the official 2014 train/val splits. The training set contains 82,783 images and the validation set (used as our test set) contains 40,504 images. There are 80 object categories. The privileged set $\mathcal{P}$ consisted of the 16 labels (20% of 80) with the lowest frequency in the training set. The remaining 64 labels formed $\bar{\mathcal{P}}$.
NUS-WIDE (chua2009nus): This dataset contains 269,648 images with 81 concept labels. We used the common split of 161,789 images for training and 107,859 for testing. The privileged set $\mathcal{P}$ consisted of the 16 labels (approx. 20% of 81) with the lowest frequency in the training set. The remaining 65 labels formed $\bar{\mathcal{P}}$.
@@ Appendix F Dataset and Preprocessing Details
Appendix G FairPO Experimental Details
G.1 Common Setup for All FairPO Variants